OpenStudy (anonymous):

the difference of two numbers is 28. their sum is 72. what is the smaller number?

OpenStudy (anonymous):

We have a system : x - y = 28 x+ y = 72 Sum them : x+x+y-y=20+72 2x = 92 x = 46 and y : 46 - y = 28 -y = 28-46

OpenStudy (anonymous):

44+28=72

OpenStudy (anonymous):

yep ! @briannamooney :please answer to my question I said in your question ;)